OpenStudy (anonymous):

Let the letters of the alphabet be A=01, B=02, . . , Z=26. For RSA encryption let n=119 and e=5. What letter is encrypted as 44? You must answer this question by "cracking" the code, that is, by decrypting rather than by trial and error. You must show or explain your work, however you accomplish obtaining an answer. (Hint: modular exponentiation may come in handy.)

myininaya (myininaya):

what do you mean what letter is encrypted as 44?

myininaya (myininaya):

\[n=p \cdot q =119\] \[e=5 \] okay so have someone's public key x is the message the encryption function is \[E(x)=x^e \mod n=c\] the decryption function is \[D(c)=c^d \mod n=x\]

myininaya (myininaya):

so what you mean the ciphertext is 44?

myininaya (myininaya):

ok and we want to decrypt this ciphertext

myininaya (myininaya):

so let's put it in the decryption function

myininaya (myininaya):

\[D(44)=44^d \mod n \]

myininaya (myininaya):

but wait we don't know d

myininaya (myininaya):

\[44^d \mod 119\] so i don't understand it says to break the code but it also says use the decryption ok we are suppose to do both

myininaya (myininaya):

we need to factor n

myininaya (myininaya):

n=119=7(17)=pq

myininaya (myininaya):

\[\phi(n)=(7-1)(17-1)=6(16)=96\]

myininaya (myininaya):

now we know e which is 5 so we can find d by solving \[de \equiv 1 \mod \phi(n)\]

myininaya (myininaya):

\[5d \equiv 1 \mod 96\]

myininaya (myininaya):

\[5d-96k=1 \text{ for some integer k }\]

myininaya (myininaya):

we need to use euclidean algorithm

myininaya (myininaya):

\[96=5(19)+1\] \[19=5(3)+4\] \[5=\] => 96-5(19)=1 =>5(-19)-96(-1)=1 so d=-19+96=77 see 5(77)-96(4)=1 okay so we have d=77 :)

myininaya (myininaya):

so back to the decryption function

myininaya (myininaya):

\[D(44)=44^d \mod n=44^{77} \mod 119\]

myininaya (myininaya):

\[44^2 \mod 119=32\] \[44^4 \mod 119 =32 \cdot 32 \mod 119=72\] \[44^8 \mod 118=72 \cdot 72 \mod 119=67\] \[44^{16} \mod 119 =67 \cdot 67 \mod 119=86\] \[44^{32} \mod 119 =86 \cdot 86 \mod 119=18\] \[44^{64} \mod 119=18 \cdot 18 \mod 119 =86\] \[44^{77} \mod 119=44^{64} \cdot 44^{13} \mod 119=(86) \cdot 44^{8} \cdot 44^{4} \cdot 44 \mod 119\]

myininaya (myininaya):

\[=86 \cdot 67 \cdot 72 \cdot 44 \mod 119\]

myininaya (myininaya):

\[=11\]

myininaya (myininaya):

and you said 11 equals k :)

myininaya (myininaya):

i see a type-0

myininaya (myininaya):

all of those are suppose to be mod 119

OpenStudy (mr.math):

I don't really know much about it, but looks like it took some effort out of you. So, you deserved the medal!

myininaya (myininaya):

lol you're so sweet

OpenStudy (turingtest):

true that, good job (I'm guessing) what is this math called?

myininaya (myininaya):

its cryptography

myininaya (myininaya):

encryption/decryption stuff network security

OpenStudy (turingtest):

right, I thought maybe there was a name for this kind of math with "modulus" in general thx

myininaya (myininaya):

number theory

OpenStudy (turingtest):

ahh...

myininaya (myininaya):

this is all number theory stuff

OpenStudy (turingtest):

thanks!

myininaya (myininaya):

np

myininaya (myininaya):

i think they even talked about RSA in my discrete math book

myininaya (myininaya):

we never talked about in class though